OpenStudy (anonymous):

Find an equation for the line with the given properties Parallel to the line 5x-y = -10 containing the points (0,0)

OpenStudy (mathmale):

Principle: parallel lines have the same slope. Action: Find the slope of the given line 5x-y=-10. I think you mean "containing the point (0,0)," (not points). Use the point-slope formula of a straight line to determine the desired equation. You know the slope of this new line and you know that this line passes through (0,0).

OpenStudy (anonymous):

apologies for the typo. Is the formula =y -y1 = m(x-x1)

OpenStudy (mathmale):

Yes, that's the point-slope formula. The given point is (0,0). What is the slope of the given line? That's the slope y ou'll want to substitute into the point-slope formula to obtain the eq'n of this new line thru (0,0).

OpenStudy (anonymous):

will try that again.

OpenStudy (mathmale):

Solve 5x-y = -10 for y. Then compare the result to y=mx+b. Identify the value of the slope, m.

OpenStudy (anonymous):

i got x=1/5y - 2

OpenStudy (anonymous):

hiya mathmale: i got x=1/5y -2

OpenStudy (anonymous):

5x-y=10 can be written as y=5x+10

OpenStudy (triciaal):

5x - y = -10 when x = 0 y = 10 when y = 0 5x -0 =-10 so x = -2 now you have points on this line find the slope

OpenStudy (triciaal):

parallel lines have the same slope so the slope for the "new" line passing through (0, 0) is 5

OpenStudy (triciaal):

eqn of line y = mx + b

OpenStudy (anonymous):

right.... got it now thank you. Thanks triciaal.

OpenStudy (mathmale):

sorry for my (unintended) absence from this discussion. I was asking you (afievans) to solve the given equation for y (not for x). 5x-y=-10 becomes y=5x+10, which has slope 5.

OpenStudy (mathmale):

Then, with slope 5 and point (0,0), the equation of the line thru this point is y-0=5(x-0), or y = 5x.

OpenStudy (mathmale):

I'm curious: How come you solved for x when asked to solve for y? Don't want to put you on the spot, but thought it important to ask you about this.

OpenStudy (anonymous):

honestly don't know how i got solving for x. guess i am just overwhelmed. maths is not my best subject...lol

OpenStudy (mathmale):

Like anything else, maths gets easier with practice!

OpenStudy (anonymous):

i know right.... thanks for helping.